OpenStudy (anonymous):

The total rainfall in Forks, Washington, this October was 8 3/8 in. In September, 60% less rain fell than in October. How much rain fell in September in Forks, Washington? Express your answer as a mixed number in simplest form.

OpenStudy (anonymous):

This one isn't multiple choice omg.

OpenStudy (anonymous):

Added the missing info :)

OpenStudy (aripotta):

ok so 60% less means 40% total. so can just do 8 3/8 times 0.4

OpenStudy (aripotta):

or\[\frac{ 67 }{ 8}\times \frac{ 4 }{ 10 }\]

OpenStudy (anonymous):

Okay, hold up. 3 7/20

OpenStudy (anonymous):

Right?

OpenStudy (aripotta):

ye

OpenStudy (anonymous):

So that's the final answer then, right?

OpenStudy (aripotta):

yep cuz that's the simplest form

OpenStudy (anonymous):

Surprised I got that one right, lol. One question left!

OpenStudy (aripotta):

yay :) i'm glad you got it right. maybe you don't need my help lol

OpenStudy (aripotta):

in your next question, post what you think the answer is first @Brookeisbubz cuz i think you'll be able to do it on your own c:

OpenStudy (anonymous):

Oh, alright! I'll do that @AriPotta